# Weighted Pseudorandom Generators for Read-Once Branching Programs via Weighted Pseudorandom Reductions

### v2

@@ 4.1 One level of the reduction
For any integer $n,k,s$ and any $\varepsilon>0$, there exists $\tau=\Omega(\varepsilon^{\frac{2}{k+1}}/\log^{2}n)$ such that there exists a $(O(k\log n),n^{O(k)},\varepsilon)$-weighted pseudorandom reduction $(\mathsf{R},w)$ from $\mathcal{P}_{(n,s)}$ to $\mathcal{P}_{(O(k\log n),s+O(\log n(\log\log n+\log 1/\tau)))}$.
@@
We invoke Lemma 4.7 with $G$ and $k$, which gives $(G^{(k)},w^{(k)})$. This WPRG fools $\mathcal{P}_{(n,s)}$ with entrywise error $\left(4\cdot\sqrt{\tau}\cdot\log n\right)^{k+1}<\varepsilon$. We define the reduction $(\mathsf{R},w):\mathsf{R}_{i}(x_{1},...,x_{r})=G^{(k)}(i,x_{1},...,x_{r}),w_{i}=w^{(k)}(i)$.
By Lemma 4.7, the weight of this reduction is bounded by $K=n^{O(k)}$. The seed length of the reduction is $\log K=O(k\log n)$, since we need $\log K=O(k\log n)$ bits to choose $i$.
Given any $f\in\mathcal{P}_{(n,s)}$, notice that each $f\circ\mathsf{R}_{i}$ is a ROBP of length $r=O(k\log n)$ and alphabet size $2^{d}=2^{s+O(\log n(\log\log n+\log 1/\tau))}$. The reason is that the transition matrix of $f\circ\mathsf{R}_{i}$ from the $j-1$-th layer to the $j$-th layer on input $x\in\{0,1\}^{d}$ can be expressed as $(f\circ\mathsf{R}_{i})^{[j-1,j]}(x)$. As $(f\circ\mathsf{R})^{[j-1,j]}(x)=f^{[n_{i,j-1},n_{i,j}]}(G(x)_{n_{i,j}-n_{i,j-1}})$ and the latter is a product of $n_{i,j}-n_{i,j-1}$ permutation matrices, the transition matrix of $f\circ\mathsf{R}_{i}$ is also a permutation matrix. Hence $f\circ\mathsf{R}_{i}$ is a permutation ROBP. Also, note that each $f\circ\mathsf{R}_{i}$ only has one accept node since its accept node is the same as that of $f$.
@@ 4.2 Recursion of reductions
We give a multi-level recursive procedure which gradually reduces the program from $\mathcal{P}_{(n,s)}$ to $\mathcal{P}_{(O(1),s+O(\log n(\log\log n+\sqrt{\log(1/\varepsilon)})))}$ with error $\varepsilon$, for any $n,s,\varepsilon$.
We start by reducing the length of the permutation ROBP to $O(\log n\sqrt{\log(1/\varepsilon)})$ and the weight to $2^{O(\log n\sqrt{\log(1/\varepsilon)})}$ with error $\varepsilon/2$. Previous works use an INW PRG to fool the reduced ROBP, which leads to an extra double logarithmic factor in the seed length. Here we use the recursion instead to avoid this extra factor.
@@
For any integer $n,s$ and any $\varepsilon>0$, there exists a $\left(O(\log n\sqrt{\log(1/\varepsilon)}),2^{\log n\sqrt{\log(1/\varepsilon)}},O(\varepsilon)\right)$-weighted pseudorandom reduction from $\mathcal{P}_{(n,s)}$ to $\mathcal{P}_{(O(1),s+O(\log n(\log\log n+\sqrt{\log(1/\varepsilon)})+\log(1/\varepsilon)))}$.
@@
Let $(\mathsf{R}^{(0)},w^{(0)}$ be the $(d_{0},W_{0},\varepsilon/2)$-weighted pseudorandom reduction from $\mathcal{P}_{(n,s)}$ to $\mathcal{P}_{(n_{1},s_{1})}$ such that $n_{1}=O(\log n\sqrt{\log(1/\varepsilon)})$ and $s_{1}=s+O(\log n(\log\log n+\sqrt{\log(1/\varepsilon)}))$. The reduction is guaranteed by Lemma 4.9 and have the following parameters:
@@
    $W_{0}=2^{O(\log n\sqrt{\log(1/\varepsilon)})}$.
Now we define a new parameter $\varepsilon^{\prime}=\varepsilon/2W_{0}$. Let
@@
For each $i=1,\ldots,l$, let $(\mathsf{R}^{(i)},w^{(i)})$ be the $(d_{i},W_{i},(\varepsilon^{\prime})^{2})$-weighted pseudorandom reduction from $\mathcal{P}_{(n_{i},s_{i})}$ to $\mathcal{P}_{(n_{i+1},s_{i+1})}$, which is guaranteed by Lemma 4.10. The reduction has the following parameters:
@@
Using Lemma 2.3, we first composite $(\mathsf{R}^{(1)},w^{(1)}),\ldots,(\mathsf{R}^{(l)},w^{(l)})$ to get a $(d^{*},W^{*},\varepsilon^{*})$-weighted pseudorandom reduction $(\mathsf{R}^{(*)},w^{(*)})$ from $\mathcal{P}_{(n_{0},s_{0})}$ to $\mathcal{P}_{(n_{l},s_{l})}$ with $\mathsf{R}^{(*)}=\mathsf{R}^{(1)}\circ\ldots\circ\mathsf{R}^{(l)}$ and $w^{(*)}=w^{(1)}\cdot\ldots\cdot w^{(l)}$. We have the following parameters:
@@
    $W^{*}=W_{1}\cdot\ldots\cdot W_{l}\leq 2^{O(l\cdot\log^{3}n_{1})}\leq 2^{O(\log^{0.1}(n/\varepsilon))}$,
@@
    $\varepsilon^{*}=\sum_{i=1}^{l}(\varepsilon^{\prime})^{2}\cdot\prod_{j=0}^{i-1}W_{j}\leq(\varepsilon^{\prime})^{2}\cdot W^{*}\cdot l\leq\varepsilon^{\prime}/2$,
@@
The last item holds because $n_{i}$ decreases rapidly. Because $n_{i-1}=2^{\sqrt[3]{n_{i}}}$, when $n_{i}\geq 2^{15}=32768$, We have $n_{i-1}\geq n^{2}_{i}$. When $n_{l}\geq 32768$, $\sum_{i=1}^{l}\frac{\log(1/\varepsilon^{\prime})}{\log n_{i}}\leq\log 1/\varepsilon^{\prime}\cdot(1/\log n_{l}+1/(2\log n_{l})+1/(4\log n_{l})+\cdots+1/(2^{l}\log n_{l}))\leq\log 1/\varepsilon^{\prime}$.
Finally, we composite $(\mathsf{R}^{(*)},w^{(*)})$ with $(\mathsf{R}^{(0)},w^{(0)})$ to get the final reduction $(\mathsf{R},w)$ from $\mathcal{P}_{(n,s)}$ to $\mathcal{P}_{(O(1),s_{l}))}$. By Lemma 2.2, the parameters are:
@@
    $W=W^{*}\cdot W_{0}\leq 2^{O(\log n\sqrt{\log(1/\varepsilon)})}$,
@@
    $\varepsilon=\varepsilon/2+\varepsilon^{\prime}/2\cdot W_{0}\leq\varepsilon$,
@@
Given this lemma, we immediately have the main theorem of this section:
@@
Let $(\mathsf{R},w)$ be the reduction from Lemma 4.11. We construct the WPRG $(G,w)$ with the same weight function $w$, and $G$ is defined as $G(x,i)=\mathsf{R}(x,i)$. By Lemma 4.11, to fool the original ROBP we only need to provide randomness for the seed of $(\mathsf{R},w)$ and also provide a random string $x$ which can fool an arbitrary ROBP in $\mathcal{P}_{(O(1),s+O(\log n(\log\log n+\sqrt{\log(1/\varepsilon)})+\log(1/\varepsilon)))}$. We can use another $\varepsilon$ error INW generator from [HPV21a] to generate this constant length $x$ and this only takes $s+O\left(\log n(\log\log n+\sqrt{\log(1/\varepsilon)})+\log(1/\varepsilon)\right)$ random bits. Also the randomness used by $(\mathsf{R},w)$ is $O(\log n\sqrt{\log(1/\varepsilon)})$. So the total seed length is as stated. The weight also directly follows from Lemma 4.11.
It is well known that a WPRG fools a Permutation ROBP with an arbitrary number of accept nodes and error $\varepsilon$ if it fools the program with only one accept node with error $\varepsilon/w$. Therefore, we have the following corollary:
@@
5 Derandomizing Regular Branching Programs
In this section, we give our improved derandomization for short-wide regular ROBPs. When $n\leq 2^{\log^{1/2}w}$ and $\varepsilon=1/\operatorname{\mathrm{poly}}(w)$, our result provides an $O(\log\frac{w}{\varepsilon})$ space derandomization, which is optimal up to a constant factor. We achieve this by constructing a new derandomization algorithm that runs in space $O(\log(w/\varepsilon)+\log n\sqrt{\log(w/\epsilon)})$, adapted from our WPRG for permutation ROBPs from the Section 4.
As a warm-up, we show that binary regular ROBPs can be transformed in logspace into binary permutation ROBPs such that they have the same acceptance probability (but probably do not compute the same function). This new transformation is given in Appendix D. After applying this transformation, we can run our WPRG of Corollary 4.13 to attain the desired derandomization for the binary case.
However for non-binary alphabets, the technique turns out to be much more complicated. In the rest of this section, we mainly focus on this case.
@@ 5.1 More preliminaries
A regular ROBP with two-way labeling is a pair $(f,\ell_{in})$ where $f$ is a regular ROBP in $\mathcal{B}_{n,s,w}$ with vertex set $V=V_{0}\cup V_{1}\cup\cdots\cup V_{n}$, edge set $E=E_{1}\cup\cdots\cup E_{n}$ and edge labeling $\ell_{out}:E\to[2^{s}]$. The function $\ell_{in}:E\to[2^{s}]$ map every edge to another label. Furthermore, for every vertex $i\in[n]$, $l_{in}$ and $l_{out}$ restricted on $E_{i}$ is a two-way labeling of the bigraph $G_{i}:=(V_{i-1},V_{i},E_{i})$.
We neglect $\ell_{out}$ in the notation since we can always consider $\ell_{out}$ to be the canonical order given by the ROBP.
@@
Let $(f,\ell_{in})$ be a regular ROBP with two-way labeling in $\mathcal{R}^{tw}_{n,s,w}$. For each $i\in[\log n]$, let $H_{t}=([2^{s}\cdot c^{t-1}],[2^{s}\cdot c^{t}],E_{H_{t}})$ be a $c$-regular bigraph with $\lambda(H_{t})\leq\lambda$. Define $\widetilde{G}_{j\to j+1}:=(V_{t-1},V_{t},E_{t})$, which is a $2^{s}$-regular bigraph with two-way labeling. Define $E(SC_{n})=\{(j,j+2^{t}):j\in[n-2^{t}],t\in[\log n],2^{t}\text{ is the largest power of 2 dividing }j\}$. For each $(j,j+2^{t})\in E(SC_{n})$ recursively define the graph $\widetilde{G}_{j\to j+2^{t}}$ as follows:
@@
Let $\{A_{i}\}_{i=1}^{n}\subset\mathbb{R}^{w\times w}$ be a sequence of doubly stochastic matrices. Let $\{B_{i,j}\}_{i,j=0}^{n}\subset\mathbb{R}^{w\times w}$ be a family of matrices such that $B_{i,j}\overset{sv}{\approx}_{\varepsilon/(10\log n)}A_{i+1}\ldots A_{j}$ Assuming that $B_{i-1,i}=A_{i}$ for all $i$. Then for any $k\in\mathbb{N}$, there exists $K=O((2n)^{k})$, $t=O(k\log n)$, a set of indices $n_{i,j}$ for each $i\in[K],j\in[t]$ with $0\leq n_{i,1}\leq\ldots\leq n_{i,t}=n$ and a set of signs $\sigma_{i}\in\{-1,0,1\}$ for each $i\in[K]$ such that:
$\sum_{i\in[K]}\sigma_{i}\cdot B_{0,n_{i,1}}B_{n_{i,1},n_{i,2}}\ldots B_{n_{i,t-1},n_{i,t}}\overset{sv}{\approx}_{\varepsilon^{k}}A_{1}A_{2}\ldots A_{n}.$
5.2 The algorithm
We give our derandomization algorithm for regular branching programs.
Theorem 5.14.

### v3

@@ 4.1 One level of the reduction
For any integer $n,k,s$ and any $\varepsilon>0$, there exists $\tau=\Omega(\varepsilon^{\frac{2}{k+1}}/\log^{2}n)$ such that there exists a $(O(k\log n),n^{O(k)},\varepsilon)$-weighted pseudorandom reduction $(\mathsf{R},\sigma)$ from $\mathcal{P}_{(n,s)}$ to $\mathcal{P}_{(O(k\log n),s+O(\log n(\log\log n+\log 1/\tau)))}$.
@@
We invoke Lemma 4.7 with $G$ and $k$, which gives $(G^{(k)},\sigma^{(k)})$. This WPRG fools $\mathcal{P}_{(n,s)}$ with entrywise error $\left(4\cdot\sqrt{\tau}\cdot\log n\right)^{k+1}<\varepsilon$. We define the reduction $(\mathsf{R},\sigma):\mathsf{R}_{i}(x_{1},...,x_{r})=G^{(k)}(i,x_{1},...,x_{r}),w_{i}=\sigma^{(k)}(i)$. By Lemma 4.7, the weight of this reduction is bounded by $K=n^{O(k)}$. The seed length of the reduction is $\log K=O(k\log n)$, since we need $\log K=O(k\log n)$ bits to choose $i$.
Given any $f\in\mathcal{P}_{(n,s)}$, notice that each $f\circ\mathsf{R}_{i}$ is a permutation ROBP of length $r=O(k\log n)$ and alphabet size $2^{d}=2^{s+O(\log n(\log\log n+\log 1/\tau))}$. The reason is that the transition matrix of $f\circ\mathsf{R}_{i}$ from the $j-1$-th layer to the $j$-th layer on input $x\in\{0,1\}^{d}$ can be expressed as $(f\circ\mathsf{R}_{i})^{[j-1,j]}(x)$. As $(f\circ\mathsf{R})^{[j-1,j]}(x)=f^{[n_{i,j-1},n_{i,j}]}(G(x)_{n_{i,j}-n_{i,j-1}})$ and the latter is a product of $n_{i,j}-n_{i,j-1}$ permutation matrices, the transition matrix of $f\circ\mathsf{R}_{i}$ is also a permutation matrix. Hence $f\circ\mathsf{R}_{i}$ is a permutation ROBP. Also, note that each $f\circ\mathsf{R}_{i}$ only has one accept node since its accept node is the same as that of $f$.
@@ 4.2 Recursion of reductions
We give a multi-level recursive procedure which gradually reduces the program from $\mathcal{P}_{(n,s)}$ to $\mathcal{P}_{(O(1),s+O(\log n(\log\log n+\sqrt{\log(1/\varepsilon)})))}$ with error $\varepsilon$.
We start by reducing the length of the permutation ROBP to $O(\log n\sqrt{\log(1/\varepsilon)})$ and the weight to $2^{O(\log n\sqrt{\log(1/\varepsilon)})}$ with error $\varepsilon/2$. Previous works use an INW PRG to fool the reduced ROBP, which leads to an extra double logarithmic factor in the seed length. Here we use iterative reductions instead to avoid this extra factor.
@@
For any integer $n,s$ and any $\varepsilon>0$, there exists a $\left(O\left(\log n\sqrt{\log(1/\varepsilon)}\right),2^{\log n\sqrt{\log(1/\varepsilon)}},O(\varepsilon)\right)$-weighted pseudorandom reduction from $\mathcal{P}_{(n,s)}$ to $\mathcal{P}_{\left(O(1),s+O\left(\log n\left(\log\log n+\sqrt{\log(1/\varepsilon)}\right)+\log(1/\varepsilon)\right)\right)}$.
@@
Let $(\mathsf{R}^{(0)},\sigma^{(0)})$ be the $(d_{0},K_{0},\varepsilon/2)$-weighted pseudorandom reduction from $\mathcal{P}_{(n,s)}$ to $\mathcal{P}_{(n_{1},s_{1})}$ such that $n_{1}=O(\log n\sqrt{\log(1/\varepsilon)})$ and $s_{1}=s+O(\log n(\log\log n+\sqrt{\log(1/\varepsilon)}))$. This reduction is guaranteed by Lemma 4.9 and have the following parameters:
@@
    $K_{0}=2^{O(\log n\sqrt{\log(1/\varepsilon)})}$.
Now we define a new parameter $\varepsilon^{\prime}=\varepsilon/2K_{0}$. Let
@@
For each $i=1,\ldots,l$, let $(\mathsf{R}^{(i)},\sigma^{(i)})$ be the $(d_{i},K_{i},(\varepsilon^{\prime})^{2})$-weighted pseudorandom reduction from $\mathcal{P}_{(n_{i},s_{i})}$ to $\mathcal{P}_{(n_{i+1},s_{i+1})}$, which is guaranteed by Lemma 4.10. The reduction has the following parameters:
@@
Using Lemma 2.3, we first composite $(\mathsf{R}^{(1)},\sigma^{(1)}),\ldots,(\mathsf{R}^{(l)},\sigma^{(l)})$ to get a $(d^{*},K^{*},\varepsilon^{*})$-weighted pseudorandom reduction $(\mathsf{R}^{(*)},\sigma^{(*)})$ from $\mathcal{P}_{(n_{0},s_{0})}$ to $\mathcal{P}_{(n_{l},s_{l})}$ with $\mathsf{R}^{(*)}=\mathsf{R}^{(1)}\circ\ldots\circ\mathsf{R}^{(l)}$ and $\sigma^{(*)}=\sigma^{(1)}\cdot\ldots\cdot\sigma^{(l)}$. We have the following parameters:
@@
    $K^{*}=K_{1}\cdot\ldots\cdot K_{l}\leq 2^{O(l\cdot\log^{3}n_{1})}\leq 2^{O(\log^{0.1}(n/\varepsilon))}$,
@@
    $\varepsilon^{*}=\sum_{i=1}^{l}(\varepsilon^{\prime})^{2}\cdot\prod_{j=0}^{i-1}K_{j}\leq(\varepsilon^{\prime})^{2}\cdot K^{*}\cdot l\leq\varepsilon^{\prime}/2$,
@@
The first three items are attained by directly plugging in our parameters. The last item holds because $n_{i}$ decreases rapidly. Note that $n_{i-1}=2^{\sqrt[3]{n_{i}}}$, and when $n_{i}\geq 2^{15}=32768$, we have $n_{i-1}\geq n^{2}_{i}$. When $n_{l}\geq 32768$, it holds that $\sum_{i=1}^{l}\frac{\log(1/\varepsilon^{\prime})}{\log n_{i}}\leq\log 1/\varepsilon^{\prime}\cdot(1/\log n_{l}+1/(2\log n_{l})+1/(4\log n_{l})+\cdots+1/(2^{l}\log n_{l}))\leq\log 1/\varepsilon^{\prime}$.
Finally, we composite $(\mathsf{R}^{(*)},\sigma^{(*)})$ with $(\mathsf{R}^{(0)},\sigma^{(0)})$ to get the final reduction $(\mathsf{R},\sigma)$ from $\mathcal{P}_{(n,s)}$ to $\mathcal{P}_{(O(1),s_{l}))}$. By Lemma 2.2, the parameters are:
@@
    $K=K^{*}\cdot K_{0}\leq 2^{O(\log n\sqrt{\log(1/\varepsilon)})}$,
@@
    $\varepsilon=\varepsilon/2+\varepsilon^{\prime}/2\cdot K_{0}\leq\varepsilon$,
@@
Now we show the main theorem of this section:
@@
Let $(\mathsf{R},\sigma)$ be the reduction from Lemma 4.11. We construct the WPRG $(G,\sigma)$ with the same weight function $\sigma$, and $G$ is defined as $G(x,i)=\mathsf{R}(x,i)$. By Lemma 4.11, to fool the original ROBP we only need to provide randomness for the seed of $(\mathsf{R},\sigma)$ and also provide a random string $x$ which can fool an arbitrary ROBP in $\mathcal{P}_{(O(1),s+O(\log n(\log\log n+\sqrt{\log(1/\varepsilon)})+\log(1/\varepsilon)))}$. We can use another $\varepsilon$ error INW generator from [HPV21a] to generate this constant length $x$ and this only takes $s+O\left(\log n\left(\log\log n+\sqrt{\log(1/\varepsilon)}\right)+\log(1/\varepsilon)\right)$ random bits. Also the randomness used by $(\mathsf{R},\sigma)$ is $O(\log n\sqrt{\log(1/\varepsilon)})$. So the total seed length is as stated. The weight also directly follows from Lemma 4.11.
It is well known that a WPRG can fool a Permutation ROBP with an arbitrary number of accept nodes and error $\varepsilon$ if it can fool the program with only one accept node with error $\varepsilon/w$. Therefore, we immediately have the following corollary:
@@
5 Derandomization for Regular Branching Programs
In this section, we show our derandomization for short-wide regular ROBPs, which indicates that when $n\leq 2^{\log^{1/2}w}$ and $\varepsilon=1/\operatorname{\mathrm{poly}}(w)$, there is a $O(\log w)$ space derandomization, where the space complexity is optimal up to constant factors. We achieve this by constructing a more general derandomization algorithm as the following.
Theorem 5.1.
There exists an algorithm that for any input regular ROBP $f$ of width $w$, length $n$, alphabet $\{0,1\}^{s}$, and any input parameter $\varepsilon>0$ as input, outputs an approximation for $\mathbb{E}[f]$ with addtive error $\varepsilon$. The algorithm uses $O(s+\log n(\log\log n+\sqrt{\log(w/\varepsilon)})+\log(w/\varepsilon))$ bits of space.
Notice that our main theorem Theorem 1.10 is a direct corollary of Theorem 5.1. We remark that one may want to further improve the space dependence on $\varepsilon$ by first using Theorem 5.1 and then calling the Richardson iteration method of [AKM+20]. This can be done, but our main purpose here is to get a derandomization in $\mathbf{L}$ exactly for short-wide ROBPs. If applying that additional operation, then the algorithm is not in $\mathbf{L}$ and also not in polynomial time.
Next as a warm-up, we show that binary regular ROBPs can be transformed in logspace to binary permutation ROBPs such that they have the same acceptance probability (but probably do not compute the same function). We are unaware of any previous work stating this transformation.
Lemma 5.2.
Proof.
$P^{\prime}$ has the same vertex set as that of $P$. The algorithm visits each layer $i$ of $P$ in the order $i=0,1,2,\ldots,n$. For layer $i$, consider the subgraph $S$ between $V_{i}$ and $V_{i+1}$. The algorithm visits each node $v\in V_{i}$ in order. For each $v\in V_{i}$, it tests if $v$ can be reached by any previously visited node of $V_{i}$ in $S$. This can be done by revisiting each previous node $v^{\prime}$ and go through the cycle that $v^{\prime}$ is in. If none of the previous node can reach $v$, then do the following traverse in $S$. Starting from $v$, choose an edge with label 00, and then go to the corresponding neighbor. For each next node $u$ visited, denote $\ell$ as the label of the adjacent edge through which we reach $u$. If the other adjacent edge $e^{\prime}$ of $u$ has a label $\ell^{\prime}$ such that $\ell^{\prime}=\ell$, then flip the label $\ell^{\prime}$. After this we traverse through $e^{\prime}$ to go to the next neighbor. Go on doing this until it reaches $v$ again.
Notice that starting from $v$ the traverse can go through every node of $V_{i}\cup V_{i+1}$ reachable from $v$. Because for each next neighbor $u$, it can only be either $v$ or a node which is not touched before in this traverse. In other words, this traverse is going through the cycle of $S$ containing $v$. Also note that after this traverse, every node reachable from $v$, because of the flipping, can have different labels for its two adjacent edges. So after we visited every $v\in V_{i}$, every node of $S$ reachable by nodes in $V_{i}$ has its two adjacent edges with different labels because of the flips. Hence the graph with these new labels is a permutation.
The traverse can be done in logspace since recording the index of a layer takes space $O(\log n)$, recording the index of a node in a layer takes space $O(\log w)$, and going through a cycle of $S$ takes space $O(\log w)$.
After this transformation, we can run our WPRG of Corollary 4.13 to attain the desired derandomization for the binary case.
However for alphabets with size larger than $2$, the problem turns out to be more complicated. In the rest of this section, we focus on the case with large alphabets.
@@ 5.1 More preliminaries
A regular ROBP with two-way labeling is a regular ROBP $f$ and every edge in $f$ has two labels, the ’outgoing label’ and the ’incoming label’. For every vertex $u$ in $V(f)\backslash V_{n}$, the outgoing labels of the outgoing edges of $u$ are distinct. For every vertex $v$ in $V(f)\backslash V_{0}$, the incoming labels of the incoming edges of $v$ are distinct. We compute $f$ according to the outgoing labels of the edges, i.e., for every $(x_{1},\ldots,x_{n})\in(\{0,1\}^{s})^{n}$, we find the unique path $(v_{0}=v_{start},v_{1},\ldots,v_{n})$ in $f$ such that for every $i\in[n]$, the edge $(v_{i-1},v_{i})$ has the outgoing label $x_{i}$, and we output 1 iff $v_{n}$ is an accepting vertex.
@@
Like bigraphs, we define $\mathsf{Rot}_{f}:([w]\times[n])\times\{0,1\}^{s}\to([w]\times[n])\times\{0,1\}^{s}$ such that $\mathsf{Rot}_{f}(u,\sigma)=(v,\sigma^{\prime})$ if there is an edge $(u,v)$ in $f$ with the outgoing label $\sigma$ and the incoming label $\sigma^{\prime}$.
@@
Let $f$ be a regular ROBP with two-way labeling in $\mathcal{R}^{tw}_{n,s,w}$. For each $i\in[\log n]$, let $H_{t}=([2^{s}\cdot c^{t-1}],[2^{s}\cdot c^{t}],E_{H_{t}})$ be a $c$-regular bigraph with $\lambda(H_{t})\leq\lambda$. Define $\widetilde{G}_{j\to j+1}:=(V_{t-1},V_{t},E_{t})$, which is a $2^{s}$-regular bigraph with two-way labeling. Define $E(SC_{n})=\{(j,j+2^{t}):j\in[n-2^{t}],t\in[\log n],2^{t}\text{ is the largest power of 2 dividing }j\}$. For each $(j,j+2^{t})\in E(SC_{n})$ recursively define the graph $\widetilde{G}_{j\to j+2^{t}}$ as follows:
@@
We use the following notation for the algorithm in Theorem 5.13 equipped with the efficiently computable spectral expanders.
Definition 5.15.
We define the algorithm $\operatorname{DerandWalk}$. Let $n,s,w\in\mathbb{N}$, and $0<\tau<\frac{1}{6\log n}$. For every $t\in[\lceil\log n\rceil]$, let $c\in\mathbb{N}$ and $H_{t}=\big{(}[2^{s}\cdot c^{t-1}],[2^{s}\cdot c^{t-1}],E_{H_{t}}\big{)}$ be a $c$-regular bigraph with a two-way labeling satisfying $\lambda(H_{t})\leq\lambda$ constructed by Lemma 5.14. For any $f\in\mathcal{R}^{tw}_{n,s,w}$, $0\leq l<r\leq n$, $u\in[w]$, $x\in\{0,1\}^{s}$ and $e_{1},e_{2}\ldots,e_{\lceil\log n\rceil}\in[c]$, we denote $\operatorname{DerandWalk}(\tau,f,l,r,u,(x,e_{1},\ldots,e_{\lceil\log n\rceil}))$ as the algorithm in Theorem 5.13 instantiated with the $l$-th layer to the $r$-th layer of $f$ and $\{H_{t}\}$, starting at node $u$ with seed $(x,e_{1},\ldots,e_{\lceil\log n\rceil})$, i.e. $\operatorname{DerandWalk}(\tau,f,l,r,u,(x,e_{1},\ldots,e_{\lceil\log n\rceil}))=\mathsf{Rot}_{\widetilde{G}_{l\to r}}(u,(x,e_{1},\ldots,e_{\log n}))$, where $\widetilde{G}_{l\to r}$ is the bigraph generated by Lemma 5.12 with the $l$-th layer to the $r$-th layer of $f$.
@@
Let $\{A_{i}\}_{i=1}^{n}\subset\mathbb{R}^{w\times w}$ be a sequence of doubly stochastic matrices. Let $\{B_{i,j}\}_{i,j=0}^{n}\subset\mathbb{R}^{w\times w}$ be a family of matrices such that $B_{i,j}\overset{sv}{\approx}_{\tau/(10\log n)}A_{i+1}\ldots A_{j}$ Assuming that $B_{i-1,i}=A_{i}$ for all $i$. Then for any $k\in\mathbb{N}$, there exists $K=O((2n)^{k})$, $t=O(k\log n)$, a set of indices $m_{i,j}$ for each $i\in[K],j\in[t]$ with $0\leq m_{i,1}\leq\ldots\leq m_{i,t}=n$ and a set of signs $\sigma_{i}\in\{-1,0,1\}$ for each $i\in[K]$ such that:
$\sum_{i\in[K]}\sigma_{i}\cdot B_{0,m_{i,1}}B_{m_{i,1},m_{i,2}}\ldots B_{m_{i,t-1},m_{i,t}}\overset{sv}{\approx}_{\tau^{k}}A_{1}A_{2}\ldots A_{n}.$
5.2 The derandomization for regular ROBPs with large alphabets
Theorem 5.17 (Restatement of Theorem 5.1).